If A =

(i 0
0 i)
, where i =
√−1, find A^2
and A^4
(Matrices)
1
Expert's answer
2020-05-28T17:59:35-0400

"A^2=\begin{pmatrix}
 i & 0 \\
 0 & i
\end{pmatrix}\begin{pmatrix}
 i & 0 \\
 0 & i
\end{pmatrix}=\begin{pmatrix}
 i*i+0*0 & i*0+0*i \\
 0*i+i*0 & 0*0+i*i
\end{pmatrix}=\begin{pmatrix}
 -1 & 0 \\
 0 & -1
\end{pmatrix}"

"A^4=\begin{pmatrix}
 -1 & 0 \\
 0 & -1
\end{pmatrix}\begin{pmatrix}
 -1 & 0 \\
 0 & -1
\end{pmatrix}=\begin{pmatrix}
 -1(-1)+0*0 & -1*0+0(-1) \\
 00(-1)-1*0 & -1(-1)+0*0
\end{pmatrix}=\begin{pmatrix}
 1 & 0 \\
 0 & 1
\end{pmatrix}"